Q111/11C Q2 (0.5/1) Q3 [1/1]CQ4(0.8/1) Q 15 11/1] Grade: 13.3/15 Print Version You measure 20 backpacks weights, and find they have a mean weight of 57 ounces. Assume the population standard deviation is 9.4 ounces. Based on this, construct a 90% confidence interval for the true population mean backpack weight. Give your answers as decimals, to two places 57 3.65 ounces
media%2Fe1b%2Fe1b7d690-44be-4dbe-b42d-8b
media%2Fe2f%2Fe2f9e90a-2e43-49ad-85a7-e1
Answer #1

1. n= 20, x = 57, σ= 9.4

At \small \alpha = 0.10, the critical value, zc = 1.645

90% confidence interval for the tre population mean

\small \bar x \pm z_c*\frac{\sigma}{\sqrt{n}}

\small 57 \pm 1.645*\frac{9.4}{\sqrt{20}}

\small \textbf{57} \pm \textbf{3.46}

2. \small \small n= 10, \ \bar x = 41, \ s = 2

At \small \alpha = 0.20 and df = 10-1=9, the critical value, tc = 1.383

Margin of error, E =

\small E= t_c*\frac{s}{\sqrt{n}} = 1.383*\frac{2}{\sqrt{10}} =\textbf{0.87}

----------------------------------------------------------------

Please repost for last question for answer.